20000$ is split into two investments one paying 5% and the other paying 6.5% to the nearest cent how much should be invested in each so the yearly intrest from the 5% investment is double the intrest from 6.5% investment

Answer by mananth(12270)

let investment at 6.5% be x
Investment at 5 % = (20000-x)
...
5%x= 2*6.5%(20000-x)
multiply by 100
5x=2*6.5(20000-x)
5x=130000-13x
18x=130000
x= 7,222.22 at6. 5%
Balance at 6.5%